OpenStudy (anonymous):

need help: If Sam can do a job in 4 days that Lisa can do in 6 days and Tom can do in 2 days, how long would the job take if Sam, Lisa, and Tom worked together to complete it

OpenStudy (anonymous):

Let t = the number of days to complete the job. t/4 + t/6 + t/2 = 1 3t + 2t + 6t = 12 11t = 12 t = 12/11 12/11 = 1 1/11 So, about 1 day and 2 hours.